Kawaiikitty:

Find the volume of the square pyramid shown. Round to the nearest whole number. The diagrams not drawn to scale . height:9cm, width:13cm a.507 cm^3 b.351 cm^3 c.38 cm^3 d.761 cm^3

dude:

The equation for volume of a square pyramid is \(V=\frac{1}{3}b_{(base~area)}\cdot h_{height}\) To find the base area multiply the width by itself, then multiply what you have times 9, finally diving it by 3 \(\large\frac{(13*13)*9}{3}\)

dude:

Yep

Kawaiikitty:

dude:

Yeah your answer was right 13*13 = 169 169*9 = 1521 1521/3 = 507

Kawaiikitty:

thank you!